OpenStudy (sneville):

Will give medal! I need help. Tom has a collection of 30 CDs and Nita has a collection of 18 CDs. Tom is adding 1 CD a month to his collection while Nita is adding 5 CDs a month to her collection. Find the number of months after which they will have the same number of CDs. A. 1 month B. 33 months C. 2 months D. 3 months

OpenStudy (jonathan34):

hello

OpenStudy (jonathan34):

T = number of CD's that Tom has. N = number of CD's that Nita has. T = 30 + x N = 18+ 5x When Tom and Nita have the same number of CDs, T = N = y 30 + x = 18 + 5x = y 12 = 4x x = 3 y = 30 + 3 = 33 Tom and Nita will each have the same number of CDs, 33, in 3 months.

OpenStudy (jonathan34):

please medal i need it !